What is the concentration of the resulting solution if 50.00 mL of a 1:5 (w/V) solution of chlorpromazine hydrochloride is diluted to 250 mL?


1
Expert's answer
2021-01-15T06:30:17-0500

"M1 = \frac{(1.5)\times10}{318.86} = 0.047 M"

for dilution

V1M1 = V2M2

"50\times0.047 = 250\times M2"

"M2 = \frac{0.047}{5} = 0.0094 M"